# From Tower to Spire: Adding the Speech Modality to a Text-Only LLM

### v1

@@ Abstract
Large language models (LLMs) have shown remarkable performance and generalization capabilities across multiple languages and tasks, making them very attractive targets for multimodality integration (e.g., images or speech). In this work, we extend an existing LLM to the speech modality via speech discretization and continued pre-training. In particular, we are interested in multilingual LLMs, such as Tower, as their pre-training setting allows us to treat discretized speech input as an additional translation language. The resulting open-source model, Spire, is able to transcribe and translate English speech input while maintaining Tower’s original performance on translation-related tasks, showcasing that discretized speech input integration as an additional language is feasible during LLM adaptation. We make our code and models available to the community.
From Tower to Spire: Adding the Speech Modality to a Text-Only LLM
Kshitij Ambilduke††thanks: Equal contribution.††thanks: Work begun during an internship at Instituto de Telecomunicações.1, Ben Peters∗2, Sonal Sannigrahi∗2,3, Anil Keshwani†4, Tsz Kin Lam5, Bruno Martins3,6, Marcely Zanon Boito††thanks: Equal contribution.7, André F.T. Martins‡2,3,8,9 1Paris-Saclay University 2Instituto de Telecomunicações 3Instituto Superior Técnico, Universidade de Lisboa 4Sapienza University of Rome 5University of Edinburgh 6INESC-ID 7NAVER LABS Europe 8ELLIS Unit Lisbon 9Unbabel Correspondence: benzurdopeters@gmail.com
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable success across various text-based natural language processing tasks Achiam et al. (2023); Touvron et al. (2023); Jiang et al. (2024); Yang et al. (2024); Alves et al. (2024); Martins et al. (2024), motivating research into extending them to other modalities. This has led to the development of multimodal LLMs (MLLMs) capable of processing speech, audio, images and video Team et al. (2023); Driess et al. (2023); Rubenstein et al. (2023); Liu et al. (2023); Tang et al. (2023); Défossez et al. (2024); Hu et al. (2024); Laurençon et al. (2024); Huang et al. (2024); Nguyen et al. (2025). For speech-LLM integration, an effortless approach is to link the output of an automatic speech recognition (ASR) system to a text-only LLM Huang et al. (2024). However, this solution fails to fully leverage the LLM’s superior language modeling capabilities for disambiguating transcripts.
@@
More popular are solutions that investigate equipping LLMs with speech processing capabilities through modality projection Shu et al. (2023); Radhakrishnan et al. (2023); Wu et al. (2023a); Tang et al. (2023); Xue et al. (2024); Hu et al. (2024). Typically, a speech foundation model generates speech representations that are mapped to the embedding space of the LLM. The speech model is then fine-tuned along with a projector on speech-to-text tasks to equip the LLM with speech processing capabilities. In this setting, key challenges include prompt overfitting and high training costs, as tuning these MLLMs requires the adaptation of the speech projector module on vast amounts of raw speech data Tang et al. (2023); Hu et al. (2024).
@@
An alternative approach for MLLMs is the use of speech discretization, where continuous speech features are transformed prior to training into sequences of “discrete speech units” (DSUs), which can be processed similarly to text Chou et al. (2023a); Zhang et al. (2023); Rubenstein et al. (2023); Chang et al. (2024); Défossez et al. (2024); Trinh et al. (2024); Maiti et al. (2024); Nguyen et al. (2025). This approach simplifies training by eliminating the need for additional parameters beyond extended embedding matrices. Finally, while both projector-based and discretization-based MLLMs have shown promising results on text-to-speech and/or speech-to-text tasks, their development has prioritized speech-centric tasks at the expense of textual performance. Currently, limited research has focused on integrating speech while preserving the LLM’s original capabilities in textual tasks Chou et al. (2023b); Huang et al. (2024).
@@
In this work we present Spire, a speech-augmented LLM built on top of the open multilingual model Tower Alves et al. (2024). Spire can process English speech and perform ASR and speech translation (ST) while maintaining Tower’s strong performance on machine translation (MT). Spire encodes speech via HuBERT-based Hsu et al. (2021) k-means clusterization, as in previous work Zhang et al. (2023); Rubenstein et al. (2023); Chang et al. (2024). We perform training in two stages: continued pre-training (CPT) and instruction tuning (IT). For the CPT stage, we use a mixture of ASR data and a small fraction of Tower’s text CPT data. For IT, we again leverage Tower’s task-specific MT data, as well as additional English ASR and ST data. Spire is trained using approximately 42.5K hours of speech. Figure 1 illustrates our training process.
We make the following contributions:
@@
We present a pipeline for integrating speech as an additional modality into an existing LLM, enabling it to transcribe and translate English speech while preserving its original MT capabilities;
@@
We compare speech integration at two stages, namely CPT and IT, demonstrating that both stages are essential for achieving optimal performance on speech tasks;
@@
We provide a reproducible pipeline to the community: all our models, datasets and scripts are made available.111https://github.com/utter-project/SpireLM
@@ 2 Related Work
An increasing number of studies have explored integrating speech into LLMs Zhang et al. (2023); Rubenstein et al. (2023); Hassid et al. (2024). For discrete speech input, Hassid et al. (2024) demonstrate the benefits of initializing a speech LLM from a text-based LLM. SpeechGPT Zhang et al. (2023) applies IT on speech-to-text cross-modal ASR, text-to-speech (TTS), and text-based question answering. AudioPALM Rubenstein et al. (2023) is trained in a multi-task fashion, similarly to SpeechGPT, but on multilingual input. Recently, VoxtLM Maiti et al. (2024) was trained jointly on DSUs and text data for ASR, TTS, and open-ended speech/text generation. Our work is most similar to that of Spirit-LM (Nguyen et al., 2025), which adapts an LLM with an interleaved mixture of DSU and text data which requires an expensive DSU-to-transcript step to create. In contrast, we adopt a more cost-effective input representation that can be extended to any language, regardless of the availability of a speech aligner. Our focus is on successfully incorporating speech input while preserving the original competence of the model, so that the resulting model can successfully perform both speech-to-text and text-only tasks. None of the aforementioned models are trained to preserve the original model’s performance in text tasks.
@@
Previous approaches involve training from scratch with task- and domain-specific data Singhal et al. (2023); Lewkowycz et al. (2022), performing CPT with a diverse training data mix designed to broadly extend the model’s knowledge Wu et al. (2023b), or doing IT use-case-specific data Chen et al. (2023). Recent work has explored combining the latter two approaches Xu et al. (2024a); Alves et al. (2024); Wei et al. (2021); Roziere et al. (2023). In our approach to integrating DSUs into Tower, we take inspiration from Alves et al. (2024) in adopting a two-step CPT+IT process. Our work differs in that we focus on adding the speech modality, whereas Alves et al. (2024) focused on increasing the multilingual capabilities of an LLM.
@@ 3 Spire: A Speech-to-Text LLM
Our goal is to equip an LLM with speech capabilities while preserving its preexisting text capabilities. As our starting point, we select Tower Alves et al. (2024) as our LLM, which was developed from Llama-2 Touvron et al. (2023) with a two-step approach: CPT on a mixture of monolingual and parallel data (TowerBase), followed by IT on translation-related tasks (TowerInstruct). We use a similar approach to extend Tower to speech. First, we perform CPT with a combination of text-only and aligned speech-to-text datasets, followed by IT using both text-only general-purpose and task-specific data curated in TowerBlocks,222https://huggingface.co/datasets/Unbabel/TowerBlocks-v0.2 alongside task-specific speech-to-text datasets. We name our model Spire.
@@ 3.1 Speech Discretization
To more easily transfer the training set-up of Tower, we use DSUs as opposed to an auxiliary speech encoder. For all speech datasets that were used, we follow recent discretization methodology Zhang et al. (2023); Rubenstein et al. (2023); Chang et al. (2024) to produce DSUs by first extracting continuous speech representations for our speech data from the 22nd layer of an English HuBERT-large model, and then using k-means clustering () to produce centroids that are used to convert our continuous speech representation into a discrete sequence of cluster IDs.333Optimizing the layer selection for feature extraction is a complex research problem Pasad et al. (2023); Mousavi et al. (2024). In this work we follow the insights from Gow-Smith et al. (2023) and Xu et al. (2024b). We train our k-means model on a collection of 235K audio files (approximately 720 hours), drawn from three speech corpora: CoVoST-2 (Wang et al., 2021b), VoxPopuli Wang et al. (2021a), and Multilingual Librispeech (MLS; Pratap et al., 2020). The CoVoST subset consists of 62K audio files from 10,049 speakers, with a maximum of 8 audio files per speaker. The VoxPopuli subset includes 65K audio files from 639 speakers, capped at 250 audio files per speaker. Finally, the MLS subset contains 107K audio files from 5,490 speakers.
@@ 3.2 SpireBase
SpireBase is trained from TowerBase-7B using both text-only and aligned speech-to-text datasets. Following previous work, we incorporate a fraction of Tower’s original training data to preserve its original performance Scialom et al. (2022); de Masson D’Autume et al. (2019).
@@ 3.2.1 Data
We use a mixture of monolingual and parallel text in Chinese (zh), Dutch (nl), English (en), French (fr), German (de), Italian (it), Korean (ko), Portuguese (pt), Russian (ru), and Spanish (es), that was sourced from the Tower training data, as well as English ASR data sourced from popular open-source ASR datasets, as reported in Table 2. Both speech and text data are downsampled to create a 6B token data mixture (5B speech; 1B text), measured by the model tokenizer.444Preliminary experiments on the data mixture led to this particular choice. Note that the 5B speech tokens include both DSUs (4.4B tokens) and their text transcriptions (0.6B tokens).
@@ 3.2.2 CPT Setup
We train SpireBase using the MegatronLLM codebase Cano et al. (2023) on 8 A100-80GB GPUs for 6 days. We use the same hyperparameters as Tower, except for the effective batch size, which in our case is 2,304. To incorporate the DSUs in the CPT stage, we extend the model’s original vocabulary by 5000 types, e.g., <extra_id_x>. This allows us to have a vocabulary that can encode both text in subword units and speech in DSUs. For the extended vocabulary, we initialize new embeddings from a multivariate Gaussian distribution. The mean of this distribution is set to the average of the original embeddings, while the covariance is derived from the empirical covariance of the original embeddings, scaled by a factor of Hewitt (2021).
@@ 3.3 SpireFull
The SpireFull model is obtained by performing instruction tuning on SpireBase using task-specific text-only and aligned speech-to-text data.
@@ 3.3.1 Data
We use a mixture of text and speech instructions for ASR, MT, and ST. The prompt formats used during training are shown in Table 1.
@@
We use TowerBlocks Alves et al. (2024), which includes high quality translation bi-texts between English and the other languages supported by Tower. It also includes instructions for the translation-related tasks of named entity recognition and automatic post-editing.
@@
In our IT set, we use 842 hours of speech across three ST training sets: FLEURS (all nine language pairs; we filter out examples whose transcriptions overlap with the FLORES devtest set), Europarl-ST Iranzo-Sánchez et al. (2020) (en {de, es, fr, it, nl, pt}), and CoVoST-2 (enzh). Since this amounts to far less data for ST than what is available for ASR, and since en{ko, ru} have only examples from the tiny FLEURS set, we augment our speech collection with pseudo-labeled data. This has been shown to be effective for other ST systems (Barrault et al., 2023). For the CV, SPGI, and GigaSpeech datasets, we select 300K ASR samples to be pseudo-labeled each. These examples are translated to all nine target languages using TowerInstruct-13B.555https://huggingface.co/Unbabel/TowerInstruct-13B-v0.1 Although this produces a very large ST corpus, not all predictions are of high quality, so we filter out examples for which the transcript-translation combination has a COMET-QE666https://huggingface.co/Unbabel/wmt22-cometkiwi-da (Rei et al., 2022b) score under 85. Finally, for each language pair, we sample 60K examples to be used in direct ST prompts and another 60K samples to be used in multi-turn prompts. This process results in 180K direct ST prompts and 180K multi-turn prompts for each language pair.777Due to our aggressive filtering, we were left with slightly fewer examples for en zh.
| ASR (CPT) | |||
| MT (CPT) | |||
| ASR (IT) | |||
| Direct ST (IT) | |||
| Multi-turn ST (IT) | |||
@@ 3.3.2 IT Training Setup
Similar to Tower, we use the chatml template OpenAI (2023) to format our instructions in dialogue form. We train models using Axolotl888https://github.com/axolotl-ai-cloud/axolotl on 4 H100-80GB GPUs for 2.7 days. We use a learning rate of and a cosine scheduler with 100 warm-up steps. We train for 4 epochs with an effective batch size of 576 and a weight decay of 0.01. We impose a maximum sequence length of 4096 and use the AdamW optimizer Loshchilov and Hutter (2019). Other hyperparameters are derived from TowerInstruct (Alves et al., 2024).

### v3

# From Tower to Spire: Adding the Speech Modality to a Translation-Specialist LLM

@@ Abstract
We introduce Spire, a speech-augmented language model (LM) capable of both translating and transcribing speech input from English into 10 other languages as well as translating text input in both language directions. Spire integrates the speech modality into an existing multilingual LM via speech discretization and continued pre-training using only K hours of speech. In particular, we adopt the pretraining framework of multilingual LMs and treat discretized speech input as an additional translation language. This approach not only equips the model with speech capabilities, but also preserves its strong text-based performance. We achieve this using significantly less data than existing speech LMs, demonstrating that discretized speech input integration as an additional language is feasible during LM adaptation. We make our code and models available to the community.
From Tower to Spire: Adding the Speech Modality to a Translation-Specialist LLM
Kshitij Ambilduke††thanks: Equal contribution. ¶Equal contribution.††thanks: Work begun during an internship at Instituto de Telecomunicações.1, Ben Peters∗††thanks: Work done while at Instituto de Telecomunicações.2, Sonal Sannigrahi∗3,4, Anil Keshwani†5, Tsz Kin Lam6, Bruno Martins2,4, André F.T. Martins¶††thanks: Work done while at Unbabel.3,4,7,8, Marcely Zanon Boito9 1ENS Paris-Saclay 2INESC-ID 3Instituto de Telecomunicações 4Instituto Superior Técnico, Universidade de Lisboa 5Sapienza University of Rome 6University of Edinburgh 7ELLIS Unit Lisbon 8TransPerfect 9NAVER LABS Europe Correspondence: benzurdopeters@gmail.com
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable success on various text-based natural language processing tasks Achiam et al. (2023); Touvron et al. (2023); Yang et al. (2024); Alves et al. (2024); Martins et al. (2024), motivating research into extending them to other modalities. This has led to the development of multimodal LMs capable of processing speech, audio, images, and video Gemini Team et al. (2023); Driess et al. (2023); Rubenstein et al. (2023); Liu et al. (2023); Tang et al. (2024); Défossez et al. (2024); Hu et al. (2024); Huang et al. (2024); Nguyen et al. (2025). However, the integration of new modalities often comes at the cost of existing capabilities Zhai et al. (2024).
@@
For speech-LLM integration, a simple approach is to link the output of an automatic speech recognition (ASR) system to a text-only LLM Huang et al. (2024). This solution, however, is prone to error propagation and depends largely on individual model quality. More popular are solutions that investigate equipping LLMs natively with speech processing capabilities through modality projection Shu et al. (2023); Radhakrishnan et al. (2023); Wu et al. (2023a); Tang et al. (2024); Xue et al. (2024); Hu et al. (2024). Typically, a speech foundation model generates speech representations that are mapped to the embedding space of the LLM, following which the model is then fine-tuned along with a projector on speech-to-text tasks to equip the LLM with speech processing capabilities. In this setting, key challenges include prompt overfitting and high training costs, as tuning these multimodal LLMs requires the adaptation of the speech projector module on vast amounts of raw speech data Tang et al. (2024); Hu et al. (2024).
@@
An alternative approach for integrating speech into a text-only LLM is to use speech discretization, where continuous speech features are transformed prior to training into sequences of “discrete speech units” (DSUs), which can be processed similarly to text Chou et al. (2023a); Zhang et al. (2023); Rubenstein et al. (2023); Chang et al. (2024); Défossez et al. (2024); Trinh et al. (2024); Maiti et al. (2024); Nguyen et al. (2025). This approach simplifies training by eliminating the need for additional parameters beyond extended embedding matrices. Finally, while both projector-based and discretization-based solutions have shown promising results on text-to-speech and speech-to-text tasks, their development has prioritized speech-centric tasks at the expense of textual performance. Furthermore, limited research has focused on integrating speech while preserving the LLM’s original capabilities in textual tasks Chou et al. (2023b); Huang et al. (2024).
@@
In this work we present Spire, a speech-augmented LLM built from the open-weight multilingual model Tower Alves et al. (2024). Spire can perform English ASR and from-English speech translation (ST) while maintaining Tower’s strong performance on machine translation (MT) across all 10 languages111en, de, fr, nl, it, es, pt, ko, ru, zh supported by Tower. Spire encodes speech via HuBERT-based Hsu et al. (2021) k-means clustering, as in previous work Zhang et al. (2023); Rubenstein et al. (2023); Chang et al. (2024). We perform training in two stages: Continued Pre-Training (CPT) and Instruction Tuning (IT). For the CPT stage, we use a mixture of ASR data and a small fraction of Tower’s text CPT data. For IT, we leverage Tower’s task-specific MT data, as well as additional English ASR and ST data. Spire is trained using only 42.5K hours of speech, differing from the large scale of data used by existing models Radford et al. (2023); Nguyen et al. (2025); Chu et al. (2024). Figure 1 illustrates our training process. We make the following contributions:
@@
We present a pipeline for integrating speech as an additional modality into an existing LLM, enabling it to transcribe and translate English speech while preserving its original text capabilities across 10 languages;
@@
We analyze speech integration at two stages, namely CPT and IT, demonstrating the necessity of both stages to achieve optimal performance across both modalities;
@@
We make our models, datasets, and scripts available to the community.222https://huggingface.co/collections/utter-project/spire-67d4253d6af8d6a0308527e0
@@ 2 Related Work
An increasing number of studies have explored integrating speech into LLMs Zhang et al. (2023); Rubenstein et al. (2023); Hassid et al. (2024). For discrete speech input, Hassid et al. (2024) demonstrate the benefits of initializing a speech LLM from a text-based LLM. SpeechGPT Zhang et al. (2023) applies IT on speech-to-text ASR, text-to-speech (TTS), and text-based question answering. AudioPALM Rubenstein et al. (2023) is trained in a multi-task fashion, similarly to SpeechGPT, but on multilingual input. Recently, VoxtLM Maiti et al. (2024) was trained jointly on DSUs and text data for ASR, TTS, and open-ended speech/text generation. Our work is most similar to SpiritLM (Nguyen et al., 2025), which adapts an LLM with an interleaved mixture of DSU and text data, which requires an expensive DSU-to-transcript step to create. In contrast, we adopt a more cost-effective input representation that can be extended to any language, regardless of the availability of a speech aligner. Our focus is on successfully incorporating speech input while preserving the original competence of the model, so that the resulting model can successfully perform both speech-to-text and text-only tasks. None of the aforementioned models are trained to preserve the original model’s performance in text tasks.
@@
Previous approaches involve training from scratch with task- and domain-specific data Singhal et al. (2023); Lewkowycz et al. (2022), performing CPT with a diverse training data mix designed to broadly extend the model’s knowledge Wu et al. (2023b), or IT on use-case-specific data Chen et al. (2023). Recent work has explored combining the latter two approaches Xu et al. (2024a); Alves et al. (2024); Wei et al. (2021); Roziere et al. (2023). In our approach to integrating DSUs into Tower, we take inspiration from Alves et al. (2024) in adopting a two-step CPT+IT process. Our work differs in that we focus on adding the speech modality, whereas Alves et al. (2024) focused on increasing the multilingual capabilities of an LLM.
@@ 3 Spire: A Speech-to-Text LLM
We introduce Spire, whose goal is to equip an LLM with speech capabilities while preserving its preexisting text capabilities. As our base LLM we choose Tower Alves et al. (2024), which was developed from Llama-2 Touvron et al. (2023) with a two-step approach: CPT on a mixture of monolingual and parallel data (TowerBase), followed by IT on translation-related tasks (TowerInstruct). We use an approach similar to Tower to extend the model to the speech modality. First, we perform CPT with a combination of text-only and aligned speech-to-text datasets, followed by IT using both text-only general-purpose and task-specific data curated in TowerBlocks,333https://huggingface.co/datasets/Unbabel/TowerBlocks-v0.2 alongside task-specific speech-to-text datasets.
We choose Tower in particular due to its competitive performance compared to other open alternatives. Tower-based models were among the best participating systems in the WMT24 general translation task Kocmi et al. (2024). Tower’s usage of open source data during the CPT phase along with the release of the TowerBlocks dataset, used in the IT phase, further motivates our choice.
@@ 3.1 Speech Discretization
To easily transfer the training setup of Tower, we use DSUs as opposed to an auxiliary speech encoder. For all speech datasets that were used, we follow recent discretization methodology Zhang et al. (2023); Rubenstein et al. (2023); Chang et al. (2024) to produce DSUs by first extracting continuous speech representations for our speech data from the 22nd layer of an HuBERT-large model, trained on 60K hours of English speech Hsu et al. (2021), and then using k-means clustering () to produce centroids that are used to convert our continuous speech representation into a discrete sequence of cluster IDs.444Optimizing the layer selection for feature extraction is a complex research problem Pasad et al. (2023); Mousavi et al. (2024). In this work we follow the insights from Gow-Smith et al. (2023) and Xu et al. (2024b). We train our k-means model on a collection of 235K audio files (approximately 720 hours), drawn from three speech corpora: CoVoST-2 (Wang et al., 2021b), VoxPopuli Wang et al. (2021a), and Multilingual LibriSpeech (MLS; Pratap et al., 2020). The CoVoST subset consists of 62K audio files from 10,049 speakers, with a maximum of 8 audio files per speaker. The VoxPopuli subset includes 65K audio files from 639 speakers, capped at 250 audio files per speaker. Finally, the MLS subset contains 107K audio files from 5,490 speakers.
@@ 3.2 SpireBase
The first CPT stage, yielding SpireBase, is trained from TowerBase-7B555We used Tower-7B models instead of the 13B or 70B versions due to its lower compute requirements. using both text-only and aligned speech-to-text datasets. Following previous work, we include a fraction of Tower’s original training data to preserve its existing performance Scialom et al. (2022); de Masson D’Autume et al. (2019).
@@ 3.2.1 Data
We use a mixture of monolingual and parallel text in Chinese (zh), Dutch (nl), English (en), French (fr), German (de), Italian (it), Korean (ko), Portuguese (pt), Russian (ru), and Spanish (es), that was sourced from the Tower training data, as well as English ASR data sourced from popular open-source ASR datasets, as reported in Table 1. Both speech and text data are downsampled to create a 6B token data mixture (5B speech; 1B text), measured by the model tokenizer.666Preliminary experiments on the data mixture led to this particular choice. Note that the 5B speech tokens include both DSUs (4.4B tokens) and their text transcriptions (0.6B tokens).
@@ 3.2.2 CPT Setup
We train SpireBase using MegatronLLM Cano et al. (2023) on 8 A100-80GB GPUs for 6 days. We use the same hyperparameters as Tower, except for the effective batch size, which in our case is 2304. To incorporate the DSUs in the CPT stage, we extend the model’s original vocabulary by 5000 types, e.g., <extra_id_x>. This allows us to have a vocabulary that can encode both text in subword units and speech in DSUs. For the extended vocabulary, we initialize new embeddings from a multivariate Gaussian distribution. The mean of this distribution is set to the average of the original embeddings, while the covariance is derived from the empirical covariance of the original embeddings, scaled by a factor of Hewitt (2021).
@@ 3.3 SpireFull
SpireFull is obtained by instruction tuning SpireBase on task-specific text and speech data.
@@ 3.3.1 Data
We use a mixture of text and speech instructions for ASR, MT, and ST. The prompt formats used during training are shown in Appendix A.2.
@@
We use TowerBlocks Alves et al. (2024), which includes high quality translation bi-texts between English and the other languages supported by Tower. It also includes instructions for the translation-related tasks of named entity recognition (NER) and automatic post-editing (APE).
@@
In our IT set, we use 842 hours of speech across three ST training sets: FLEURS (all nine language pairs; we filter out examples whose transcriptions overlap with the FLORES devtest set), Europarl-ST Iranzo-Sánchez et al. (2020) (en {de, es, fr, it, nl, pt}), and CoVoST-2 (enzh). Since this amounts to far less data for ST than ASR, and since en{ko, ru} have only examples from the tiny FLEURS set, we augment our speech collection with pseudo-labeled data, which has been effective for other ST systems (Barrault et al., 2023). We select 300k ASR examples each from CV, SPGI, and GigaSpeech and translate them to all nine target languages using TowerInstruct-13B.777https://huggingface.co/Unbabel/TowerInstruct-13B-v0.1 We then filter examples whose transcript-translation combination has a COMET-QE888https://huggingface.co/Unbabel/wmt22-cometkiwi-da (Rei et al., 2022b) score under 85. Finally, for each language pair, we sample 60K examples to be used in direct ST prompts and another 60K to be used in multi-turn prompts. This results in 180K direct ST prompts and 180K multi-turn prompts for each language pair.999Due to our aggressive filtering, we were left with slightly fewer examples for en zh. The prompt formats are shown in Appendix A.2.
@@ 3.3.2 IT Training Setup
We use the chatml template OpenAI (2023) to format our instructions in dialogue form. We train models using Axolotl101010https://github.com/axolotl-ai-cloud/axolotl on 4 H100-80GB GPUs for 2.7 days. We use a learning rate of and a cosine scheduler with 100 warm-up steps. We train for 4 epochs with an effective batch size of 576 and a weight decay of 0.01. We impose a maximum sequence length of 4096 and use the AdamW optimizer Loshchilov and Hutter (2019). Other hyperparameters are derived from TowerInstruct (Alves et al., 2024).